# Lessons learned from field demonstrations of model predictive control and reinforcement learning for residential and commercial HVAC: A review

### v1

@@ Abstract
A large body of simulation research suggests that model predictive control (MPC) and reinforcement learning (RL) for heating, ventilation, and air-conditioning (HVAC) in residential and commercial buildings could reduce energy costs, pollutant emissions, and strain on power grids. Despite this potential, neither MPC nor RL has seen widespread industry adoption. Field demonstrations could accelerate MPC and RL adoption by providing real-world data that support the business case for deployment. This paper reviews 24 field demonstrations of MPC and RL in residential buildings and 80 in commercial buildings. After presenting demographic information – such as experiment scopes, locations, and durations – this paper analyzes experiment protocols and their influence on performance estimates. We find that 71% of the reviewed field demonstrations use experiment protocols that may lead to unreliable performance estimates. Over the remaining 29% that we view as reliable, the weighted-average cost savings, weighted by experiment duration, are 16% in residential buildings and 13% in commercial buildings. While these savings are potentially attractive, making the business case for MPC and RL also requires characterizing the costs of deployment, operation, and maintenance. Only 13 of the 104 reviewed papers report these costs or discuss related challenges. Based on these observations, we recommend directions for future field research, including: Improving experiment protocols; reporting deployment, operation, and maintenance costs; designing algorithms and instrumentation to reduce these costs; controlling HVAC equipment alongside other distributed energy resources; and pursuing emerging objectives such as peak shaving, arbitraging wholesale energy prices, and providing power grid reliability services.
@@
HVAC , model predictive control , reinforcement learning , field demonstrations , residential buildings , commercial buildings
journal: Renewable and Sustainable Energy Reviews
@@
[1]organization=School of Mechanical Engineering, Purdue University, addressline=610 Purdue Mall, postcode=47907, city=West Lafayette, IN, country=United States
@@
[2]organization=Department of Civil, Environmental and Architectural Engineering, University of Colorado, addressline=$1$ Engineering Drive, postcode=80309, city=Boulder, CO, country=United States
@@
[3]organization=National Renewable Energy Laboratory, addressline=Denver West Parkway, postcode=80401, city=Golden, CO, country=United States
@@
[4]organization=Elmore Family School of Electrical and Computer Engineering, Purdue University, addressline=610 Purdue Mall, postcode=47907, city=West Lafayette, IN, country=United States
@@ 1 Introduction
Heating, ventilation, and air conditioning (HVAC) systems in residential and commercial buildings cause about 15% of global greenhouse gas emissions [^1] From [1], heating residential and commercial buildings accounts for 10% of global greenhouse gas emissions. From [2] and [3], cooling contributes 5% (3.2% from energy use and 1.8% from refrigerants). and cost on the order of \$1 trillion per year in energy bills [^2] From [4], residential and commercial buildings use about 32% of global annual final energy, of which HVAC systems use about 57% (or about 18% of the global total). From [5], global annual final energy expenditures are about \$10 trillion (inflation-adjusted to 2024 dollars). A rough estimate of annual global expenditures on HVAC energy is therefore \$1.8 trillion.. Over the last two decades, researchers have shown that advanced control methods – such as model predictive control (MPC) and reinforcement learning (RL) – can significantly mitigate greenhouse gas emissions, reduce energy costs, and turn HVAC systems into active participants in the power grid. Despite these research efforts, however, HVAC industry adoption of MPC and RL remains slow due to market barriers, deployment challenges, and other factors [6]. This paper aims to move the technology closer to real-world adoption at scale by (1) critically reviewing what the research community has learned from field demonstrations of MPC and RL and (2) highlighting important directions for future work.
@@
Advanced HVAC control research has at least 35 years of history. Early studies focused on open-loop optimal control systems, which plan future control actions and implement them without feedback from real-time measurements or updated forecasts [7, 8]. Open-loop optimal control concepts were later extended to the framework of MPC [9]. MPC plans control actions over a receding forecast horizon, executes the first planned action, updates measurements and forecasts, then repeats the process at the next time step. While MPC makes use of predictive models to optimize control actions, RL is a potentially model-free alternative, initially demonstrated through an action-reward framework for HVAC systems [10]. Rather than explicitly using a mathematical model, RL can learn a control policy from empirical data or from model-based simulation data, making it potentially suitable for applications where developing models is costly or impractical. In a parallel line of work, data-enabled predictive control (DeePC) has evolved as a model-free variant of MPC [11]. DeePC applies Willems’ Fundamental Lemma [12] of behavioral systems theory to predict system responses directly from data, eliminating the need for an explicit system model. This makes DeePC potentially attractive for HVAC systems with complex and dynamic characteristics that make mathematical modeling difficult.
@@
Research on MPC and RL for HVAC control has grown significantly in recent years, yet experimental research remains limited. Fig. 1 shows the cumulative growth in field and non-field papers on MPC and RL since 1997. Only 4% of 2,892 relevant papers focus on field demonstrations. This limited focus is most likely due to the costs, time constraints, and other challenges associated with field demonstrations, such as gaining data access and control authority in real buildings. Despite these obstacles, field demonstrations could show the practical benefits of MPC and RL to key stakeholders, such as decision-makers in HVAC companies and government funding agencies [13]. Field demonstrations also help identify unforeseen implementation challenges and improve our understanding of how advanced HVAC control systems handle the real-world uncertainties that simulation studies often overlook [6].
@@
Despite the importance of field demonstrations in advancing real-world adoption, the research community currently lacks a systematic and comprehensive review of field demonstrations. Sturzenegger et al. [14] and Blum et al. [15] proposed and tested their own MPC systems in real-world commercial buildings while reviewing a limited number of publications on field demonstrations of MPC in commercial buildings, covering 10 and 17 studies, respectively. Similarly, Pergantis et al. [16] conducted a partial review of field demonstrations of MPC and RL in residential buildings, analyzing 12 papers, and deployed their own MPC system in a real house. This paper’s authors recently reviewed field demonstrations of MPC and RL in commercial buildings, examining 56 field tests reported in 36 peer-reviewed publications between 2005 and 2024 [17]. Saloux [18] conducted a broader partial review, focusing on methods to estimate energy savings from field implementations and referencing approximately 40 papers spanning residential and commercial buildings. Section 2 of this paper discusses the five studies mentioned above, along with other relevant literature reviews. Building on past efforts, this paper provides the most comprehensive review to date, analyzing all peer-reviewed field demonstrations of which the authors are aware. In addition to reviewing previous studies, this paper looks at emerging trends and patterns in field demonstrations and proposes experiment protocols and guidelines to support reliable and scalable implementation of advanced HVAC control.
@@ 1.2 Organization of this paper
This paper is structured as follows: Section 2 positions this paper relative to other literature reviews on general building control strategies, MPC, RL, and field demonstrations. Section 3 describes the methodology, including the search strategy, savings evaluation approach, and statistical methods used in the study. Section 4 presents demographic information, including locations, weather conditions, and experiment scopes. Section 5 highlights several important experiment protocol choices – including benchmarking methods, controlled building portions, and test durations – and critically evaluates the reliability of existing experimental work. Section 6 discusses electricity rate structures, building properties, control architectures, weather, and implementation challenges. This section also provides recommendations for future field studies and highlights limitations that future literature reviews could address. Section 7 concludes the paper. The appendix contains two summary tables that present an overview of all the residential (LABEL:tab:res) and commercial (LABEL:tab:com) field demonstrations reviewed here.
@@ 2 Related literature reviews
In recent years, a growing number of literature review papers have analyzed the application of advanced control – including MPC, RL, and other methods – to HVAC systems in residential and commercial buildings. This section summarizes these reviews, categorizing them according to their focus. To establish context, we present a brief discussion of literature reviews on general building control methods before delving into those that specifically address MPC and RL.
2.1 General building control reviews
@@
Early reviews of building control strategies focused primarily on optimal and supervisory HVAC control. Wang and Ma [21] reviewed optimal and supervisory HVAC control methods, early model-based approaches, and broader classifications of advanced building controls. Dounis and Caraiscos [22] reviewed advanced building controls, focusing on multi-agent systems and computational intelligence, including MPC, for energy efficiency and occupant comfort. In [23], Naidu and Rieger reviewed ‘hard’ control approaches, including MPC and robust control, while in [24], they reviewed ‘soft’ control techniques, including fuzzy logic and neural networks, as well as hybrids of ‘hard’ and ‘soft’ control. A detailed analysis of intelligent control strategies, particularly neuro-fuzzy and genetic-fuzzy approaches, is provided by Mirinejad et al. [25]. Shaikh et al. [26] reviewed optimal control systems for energy and comfort management in smart sustainable buildings, highlighting a variety of control methods, such as MPC and learning-based methods. A review of building control methods was carried out, from conventional techniques such as proportional-integral-derivative control to emerging approaches such as fuzzy logic and neural networks by Royapoor et al. [27], focusing on challenges in adopting RL and integrating renewable energy systems.
Recent reviews have highlighted occupant-centric control, which integrates real-time occupancy data and occupant comfort feedback into adaptive control strategies. Jung and Jazizadeh [28] discussed human-in-the-loop HVAC operations, pointing out occupancy detection, comfort profiling, and data-driven HVAC optimization. Naylor et al. [29] considered occupant-centric approaches to real-time energy management and demand response. Park et al. [30] conducted a critical review of occupant-centric control field implementations, identifying practical challenges in deploying occupant-driven control strategies in real-world settings. Veselý and Zeiler [31] studied personalized conditioning systems and their impact on thermal comfort and energy efficiency. They found that localized HVAC control can reduce overall energy consumption while improving individual comfort. Furthermore, the integration of occupant behavior and dynamic building requirements into control systems is highlighted by Liu et al. [32], which reviews advanced controls that emphasize energy reliability, flexibility, and occupant-centric systems.
Within the fields of advanced design and control for building applications, numerical optimization methods often play a central role. Past literature review papers have explored optimization methods, such as those of Evins [33] for sustainable building design, Lazos et al. [34] for energy management in commercial buildings with weather forecasting inputs, and Wang et al. [35] for multi-criteria decision-making for renewable energy systems. Frameworks for optimizing building automation and control systems, such as the one presented by Aste et al. [36], focus on bridging the performance gap and improving decision-making under uncertainty. Further investigations have been carried out on computational optimization strategies by Banos et al. [37], demand-side management by Esther and Kumar [38], and multi-objective frameworks to balance energy efficiency, thermal comfort, and indoor air quality by Al Mindeel et al. [39]. Pallonetto et al. [40] also examined the significance of optimization in demand response schemes in homes.
@@ 2.2 Reviews of MPC for buildings
Review papers on MPC in HVAC systems can be grouped into three categories: (1) those that provide a comprehensive overview of the MPC approach, (2) those that focus on specific applications or aspects of building control, and (3) those that discuss MPC as part of an overall review of building control. In the first category, Henze [41] provided an early perspective on MPC capabilities in buildings, revealing its potential to significantly impact energy management and optimize intelligent building operations. Afram and Janabi-Sharifi [42] published one of the most complete early literature reviews on MPC, with a systematic framework for its implementation in HVAC systems, including detailed classifications, performance factors, and practical applications. Killian and Kozek [43] addressed ten questions about MPC implementation in building control, studying its challenges, potential, and future prospects, with a focus on the practical steps required for real-world application. Serale et al. [44] developed a framework for MPC in buildings, presenting a common taxonomy and addressing challenges in defining control problems, while demonstrating its potential to improve energy efficiency in HVAC systems. Expanding on these themes, Drgoňa et al. [45] offered the most comprehensive review of MPC for buildings, including a unified framework that brings together theoretical foundations, strategies for practical implementation, and performance assessment methods to close the gap between research and real-world applications. Yao and Shekar [46] provided a deeper look at MPC for HVAC systems, classifying its different types, and addressing important issues such as plant-model mismatches and disturbances. The review also highlighted optimization methods and practical approaches for implementation. In a recent study, Taheri et al. [47] presented an extensive review of MPC for HVAC systems, addressing topics such as prediction horizons and constraints, looking at several MPC methods, and recommending directions for future research.
While the aforementioned reviews focused solely on MPC, another group of literature reviews investigated MPC within broader frameworks or with other control techniques. In this context, Benndorf et al. [48] incorporated MPC into their broader review of energy performance optimization in buildings, discussing its potential within a framework that also deals with semantic interoperability and fault detection. Similarly, Mariano-Hernandez et al. [49] studied energy management strategies, such as MPC and static optimization, with applications to demand-side management and fault detection. With an emphasis on data-driven predictive control approaches, Kathirgamanathan et al. [50] highlighted how these methods can be used to unlock building energy flexibility through advanced modeling techniques, grid integration, and practical applications. Although not strictly a review, Stoffel et al. [51] compared MPC variants (white-box, gray-box, black-box, and approximate MPC) with RL, evaluating energy savings, comfort, and computational effort through simulations. Model-based and model-free predictive control strategies were the focus of a review by Xin et al. [52], exploring their applications in cooling, heating, and integrated systems while highlighting strengths, limitations, and advancements in methods such as RL.
@@
Beyond these comprehensive and broader reviews, others focused on targeted aspects of MPC applications, addressing specific challenges in HVAC and building control. For instance, with a focus on commercial buildings, Hilliard et al. [53] examined trends and opportunities for MPC, emphasizing its benefits in using building thermal mass to optimize energy use and occupant comfort. Similarly, Rockett and Hathway [54] critically evaluated MPC’s practical implementation challenges, highlighting the role of predictive modeling and sensor integration to achieve energy savings and operational efficiency. Shifting the focus to occupant-centric control, Mirakhorli and Dong [55] focused on MPC’s potential to improve HVAC energy efficiency and comfort through advanced modeling, prediction techniques, and field studies. Thieblemont et al. [56] examined MPC applications for thermal energy storage in buildings, emphasizing the use of weather data forecasts to optimize heating and cooling operations. The role of MPC in renewable energy systems was also covered by Tarragona et al. [57], who emphasized how MPC can be used to manage variability in renewable energy sources and optimize hybrid energy storage. Afram et al. [58] highlighted the ability of neural-network-based MPC to handle nonlinear HVAC dynamics, demonstrating its energy optimization potential through a residential case study.

### v6

@@ Abstract
A large body of simulation research suggests that model predictive control (MPC) and reinforcement learning (RL) for heating, ventilation, and air-conditioning (HVAC) in residential and commercial buildings could reduce energy costs, pollutant emissions, and strain on power grids. Despite this potential, neither MPC nor RL has seen widespread industry adoption. Field demonstrations could accelerate MPC and RL adoption by providing real-world data that support the business case for deployment. Here we review 24 papers that document field demonstrations of MPC and RL in residential buildings and 80 in commercial buildings. After presenting demographic information – such as experiment scopes, locations, and durations – this paper analyzes experiment protocols and their influence on performance estimates. We find that 71% of the reviewed field demonstrations use experiment protocols that may lead to unreliable performance estimates. Over the remaining 29% that we view as reliable, the weighted-average cost savings, weighted by experiment duration, are 16% in residential buildings and 13% in commercial buildings. While these savings are potentially attractive, making the business case for MPC and RL also requires characterizing the costs of deployment, operation, and maintenance. Only 13 of the 104 reviewed papers report these costs or discuss related challenges. Based on these observations, we recommend directions for future field research, including: Improving experiment protocols; reporting deployment, operation, and maintenance costs; designing algorithms and instrumentation to reduce these costs; controlling HVAC equipment alongside other distributed energy resources; and pursuing emerging objectives such as peak shaving, arbitraging wholesale energy prices, and providing power grid reliability services.
@@
HVAC , buildings , model predictive control , reinforcement learning , field demonstrations
@@
[1]organization=School of Mechanical Engineering, Purdue University, city=West Lafayette, IN, country=USA
@@
[2]organization=Department of Civil, Environmental and Architectural Engineering, University of Colorado, city=Boulder, CO, country=USA
@@
[3]organization=National Renewable Energy Laboratory, city=Golden, CO, country=USA
@@
[4]organization=Elmore Family School of Electrical and Computer Engineering, Purdue University, city=West Lafayette, IN, country=USA
[…]
[5]organization=Trane Technologies, Residential R&D Group, city=Tyler, TX, country=USA
{graphicalabstract}[Uncaptioned image]
@@ 1 Introduction
Heating, ventilation, and air conditioning (HVAC) systems in residential and commercial buildings cause about 15% of global greenhouse gas emissions [^1] From [1], heating residential and commercial buildings accounts for 10% of global greenhouse gas emissions. From [2] and [3], cooling contributes 5% (3.2% from energy use and 1.8% from refrigerants). and cost on the order of \$1 trillion per year in energy bills [^2] From [4], residential and commercial buildings use about 32% of global annual final energy, of which HVAC systems use about 57% (or about 18% of the global total). From [5], global annual final energy expenditures are about \$10 trillion (inflation-adjusted to 2024 dollars). A rough estimate of annual global expenditures on HVAC energy is therefore \$1.8 trillion.. Over the last two decades, researchers have shown that advanced control methods – such as model predictive control (MPC) and reinforcement learning (RL) – can significantly mitigate greenhouse gas emissions, reduce energy costs, and turn HVAC systems into active participants in the power grid. Despite these research efforts, however, HVAC industry adoption of MPC and RL remains slow due to market barriers, deployment challenges, and other factors [6, 7]. This paper aims to move the technology closer to real-world adoption at scale by (1) critically reviewing what the research community has learned from field demonstrations of MPC and RL and (2) highlighting important directions for future work.
@@
Advanced HVAC control research has at least 35 years of history. In the 1990s and early 2000s, studies mainly focused on open-loop optimal control systems, which plan future control actions and implement them without feedback from real-time measurements or updated forecasts [8, 9]. Open-loop optimal control concepts were later extended to the framework of MPC [10]. MPC plans control actions over a receding forecast horizon, executes the first planned action, updates measurements and forecasts, then repeats the process at the next time step. While MPC uses predictive models to optimize control actions, RL is a potentially model-free alternative that was first implemented for HVAC systems in [11, 12]. Instead of directly using a mathematical model, RL can learn a control policy from real-world data or data from model-based simulations. This means that RL may be useful in cases in which developing models would be costly or impractical. In a parallel line of work, data-enabled predictive control (DeePC) has emerged as a model-free variant of MPC [13]. DeePC applies Willems’ Fundamental Lemma [14] of behavioral systems theory to predict system responses directly from data, eliminating the challenge of developing an explicit system model. This makes DeePC potentially attractive for HVAC systems with complex and dynamic characteristics that make mathematical modeling difficult.
@@
Research on MPC and RL for HVAC control has grown significantly in recent years, but experimental research remains limited. Fig. 1 shows the cumulative growth in field and non-field papers on MPC and RL since 1997333To create this figure, we used a comprehensive Scopus search with a wide array of keywords related to MPC, RL, HVAC, and buildings. The period covered for MPC begins in 1997 with the first journal paper on MPC [10]. For RL, it begins in 2003 with the first RL paper [12]. The ending year is the time of submission of this paper.. Only 4% of 2,892 relevant papers focus on field demonstrations. This limited focus is most likely due to the costs, time constraints, and other challenges associated with field demonstrations, such as gaining data access and control authority in real buildings. Despite these obstacles, field demonstrations could show the practical benefits of MPC and RL to key stakeholders, such as decision-makers in HVAC companies and government funding agencies [15]. Field demonstrations also help identify unforeseen implementation challenges and improve our understanding of how advanced HVAC control systems handle the real-world uncertainties that simulation studies often overlook [7].
@@
Despite the importance of field demonstrations in advancing real-world adoption, the research community currently lacks a systematic and comprehensive review of field demonstrations. Previous efforts have provided limited coverage. For example, Sturzenegger et al. [16] and Blum et al. [17] reviewed small numbers of papers on field demonstrations of MPC in commercial buildings, covering 10 and 17 papers, respectively. Sturzenegger et al. and Blum et al. provided these partial reviews as context for their own field developments of MPC systems. This paper’s authors recently reviewed field demonstrations of MPC and RL in commercial buildings, covering 56 field tests from 36 papers from 2005 to 2024, and reported ongoing deployment challenges [6]. Pergantis et al. [18] surveyed 12 papers on field demonstrations of MPC and RL in residential buildings in a paper that also developed and field-tested their own MPC system. More recently, Saloux [19] carried out a broader partial review of MPC field demonstrations, focusing on methods to estimate energy savings from field implementations and referencing about 40 papers on residential and commercial buildings. Section 2 delves deeper into these five studies, as well as other relevant review papers on general HVAC control strategies, MPC, and RL. Building on past efforts, this paper provides the most comprehensive review to date, evaluating all peer-reviewed field demonstrations of which the authors are aware. In addition to reviewing previous publications, this paper looks at emerging trends and patterns in field demonstrations and proposes experiment protocols and guidelines to support reliable and scalable implementation of advanced HVAC control.
@@ 1.2 Organization of this paper
This paper is structured as follows. Section 2 compares this paper to previous review papers on general HVAC control strategies, MPC, RL, and field demonstrations. Section 3 discusses the methodology, which includes the search strategy, savings evaluation approach, and statistical methods used. Section 4 presents demographic information, including locations, weather conditions, and experiment scopes. Section 5 highlights several important experiment protocol choices – including benchmarking methods, controlled building portions, and test durations – and critically evaluates the reliability of existing experimental work. Section 6 discusses electricity rate structures, building properties, control architectures, weather, and implementation challenges. This section also provides recommendations for future field studies and highlights limitations that future literature reviews could address. Section 7 concludes the paper. A summarizes the nomenclature used in this paper. B contains two summary tables that present an overview of all the residential (LABEL:tab:res) and commercial (LABEL:tab:com) field demonstrations reviewed here.
@@ 2 Related literature reviews
In recent years, a growing number of literature review papers have analyzed the application of advanced control – including MPC, RL, and other methods – to HVAC systems in residential and commercial buildings. This section summarizes these literature review papers and categorizes them according to their focus. To establish context, we begin with a brief discussion of literature reviews on general HVAC control strategies before delving into those that specifically address MPC, RL, and field demonstrations.
2.1 General HVAC control reviews
@@
Early reviews on HVAC control focused primarily on optimal and supervisory control. Wang and Ma [21] reviewed optimal and supervisory HVAC control, early model-based approaches, and broader classifications of advanced building controls. With a particular focus on multi-agent systems, Dounis and Caraiscos [22] surveyed advanced building controls for energy efficiency and occupant comfort. In [23], Naidu and Rieger reviewed ‘hard’ control approaches, including proportional-integral-derivative (PID), optimal, model predictive, robust, nonlinear, and adaptive, while in [24], they reviewed ‘soft’ control techniques, including artificial neural networks (ANNs), fuzzy logic, genetic algorithms, as well as hybrids of ‘hard’ and ‘soft’ control. Mirinejad et al. [25] reviewed intelligent control strategies, with a particular focus on neuro-fuzzy and genetic-fuzzy approaches. In a comprehensive review of optimal control systems for energy and comfort management in smart buildings, Shaikh et al. [26] highlighted a variety of control methods, energy and comfort related trends, and future directions. Royapoor et al. [27] reviewed building control methods, ranging from conventional proportional-integral-derivative control to emerging methods like fuzzy logic and ANNs. Their review addressed challenges in adopting RL and integrating renewable electricity generation into power systems.
Occupant-centric control (OCC) has drawn recent attention in the context of building control. OCC integrates real-time occupancy data and occupant comfort feedback into control algorithms. Naylor et al. [28] provided a structured survey of OCC strategies and focused on the relevant sensing technologies and categorizing implementation approaches into different groups. Jung and Jazizadeh [29] conducted a performance-based analysis of OCC systems and provided a taxonomy to evaluate occupancy- and comfort-driven approaches. Focusing on field implementations, Park et al. [30] reviewed practical challenges in deploying OCC strategies and discussed future directions. According to a review by Liu et al. [31], OCC is a transformative strategy which includes occupant presence, preferences, and interactions to enhance HVAC operation for energy efficiency and thermal comfort.
Other review papers have focused on optimization-based strategies in order to improve system performance and decision-making. An early review by Wang et al. [32] focused on multi-criteria decision-making approaches to assess renewables across different sustainability dimensions. Banos et al. [33] reviewed computational optimization strategies applied to different energy sectors. According to a review by Evins [34], computational optimization, particularly multi-objective methods, improves sustainable building design in envelopes, building control, and renewables. Focusing on commercial buildings, Lazos et al. [35] reviewed the role of weather forecasts in optimal operation of energy management systems, focusing on load and generation predictions, and outlined an integrated framework for better control and energy savings. Focusing on the residential demand-side management, Esther and Kumar [36] presented a comprehensive review of various optimization approaches, such as deterministic, stochastic, and game-theoretic models. In addition, Aste et al. [37] presented a framework for optimal building automation, which aims to improve decision-making under uncertainty and close the performance gap. More recently, 82 multi-objective studies were covered by Al Mindeel et al. [38], targeting energy savings, thermal comfort, and indoor air quality.
@@ 2.2 Reviews of MPC for buildings
Since the early 21st century, research on MPC for buildings has grown, with authors performing reviews with different scopes, some focusing only on MPC while others addressed it alongside other advanced control techniques. Henze [39] provided an early perspective on MPC capabilities in buildings and revealed its potential to significantly impact energy management and optimize intelligent building operations. One of the most comprehensive early reviews on MPC is that of Afram and Janabi-Sharifi [40], which presented a systematic framework for MPC implementation in HVAC systems, from classifications and performance factors to practical applications. Killian and Kozek [41] posed ten questions about MPC implementation in building control, studying its challenges, potential, and future prospects, with a focus on practical steps for real-world application. Serale et al. [42] developed a detailed framework for MPC in buildings and presented a common taxonomy by addressing challenges in defining control problems, while demonstrating its potential to improve energy efficiency. Drgoňa et al. [43] offered the most comprehensive review of MPC for buildings to date, including a unified framework that brings together theoretical foundations, practical implementation strategies, and performance assessment methods to close the gap between research and real-world applications. Another derailed review was given by Yao and Shekar [44], classifying MPC types, addressing plant-model mismatches and disturbances, and highlighting optimization and implementation strategies. A recent review by Taheri et al. [45] addressed topics such as prediction horizons and constraints, looking at several MPC methods and future research directions.
@@
Some topical reviews look at the targeted aspects of MPC applications, addressing specific challenges in building control. On commercial buildings, Hilliard et al. [46] examined trends and opportunities for MPC and highlighted its benefits while using building thermal mass to optimize energy use and occupant comfort. On a similar subject, Rockett and Hathway [47] analyzed the challenges of MPC implementation in commercial buildings, pointing out the importance of predictive modeling and sensor integration for energy and operational efficiency. On occupancy behavior based MPC, Mirakhorli and Dong [48] investigated the potential of MPC to improve energy efficiency and occupant comfort via advanced modeling, prediction techniques, and field studies. Afram et al. [49] showed that using ANN-based MPC could help address non-linear HVAC dynamics and improve energy efficiency, as demonstrated through a residential case study. Thieblemont et al. [50] reviewed MPC applications for thermal energy storage in buildings and highlighted how weather forecasts were used to optimize heating and cooling operations. Lastly, Tarragona et al. [51] investigated the role of MPC in renewable energy sources, in addition to building-level controls. They also stressed the use of MPC in managing renewable energy variability and operating hybrid energy storage systems.
Some researchers also reviewed MPC within broader frameworks of building controls. Benndorf et al. [52] incorporated MPC in their review of energy performance optimization in buildings and discussed its potential within a framework that deals with semantic interoperability and fault detection. Similarly, Mariano-Hernandez et al. [53] reviewed energy management strategies, in particular MPC and static optimization, for demand-side management and fault detection. Kathirgamanathan et al. [54] reviewed the role of data-driven predictive controls in unlocking building energy flexibility through advanced modeling, grid integration, and practical applications. While not strictly a review, Stoffel et al. [55] assessed energy savings, comfort, and computational effort through simulations and compared MPC variants, including white-box, gray-box, black-box, and approximate MPC, to RL. Xin et al. [56] focused on model-based and model-free predictive controls and highlighted their applications in cooling, heating, and integrated systems while highlighting strengths, limitations, and advancements in methods like RL.